# From Ambiguity to Verdict: A Semiotic-Grounded Multi-Perspective Agent for LLM Logical Reasoning

### v1

@@ B.1 Formal Syntax and Validation of FOL
FOL CFG Grammar. To ensure the well-formedness of FOL expressions, we implement a symbolic parser using the nltk library Bird (2006). Specifically, we define a context-free grammar (CFG) to support automatic parsing and validation of logical formulas throughout our pipeline:
@@ C.1 Benchmarks
RepublicQA is a philosophical reasoning benchmark derived from Plato’s Republic, featuring complex logical reasoning problems that require models to evaluate the correctness of philosophical statements given contextual premises. RepublicQA aims to assess models’ capabilities in handling abstract philosophical concepts, moral reasoning, and classical logical arguments while maintaining high complexity in both language and reasoning patterns. The benchmark is carefully curated to reflect fundamental questions in western philosophy. Each example requires models to engage with sophisticated philosophical arguments, often involving conditional reasoning, moral principles, and abstract conceptual relationships. To ensure logical consistency, examples are structured around classical philosophical dialogues and arguments, requiring multi-step reasoning to determine whether conclusions follow from the given premises about justice, virtue, and political philosophy.
Figure 7: Answer distribution across different benchmarks.
@@ C.2 Baselines
Naive Prompting.
This approach involves directly presenting the question to the model without any additional instructions or intermediate reasoning steps. The model is expected to produce the final answer based solely on its pretrained knowledge. It does not encourage or structure the reasoning process, making it suitable only for simple or factual queries.
Chain-of-Thought (CoT).
Chain-of-Thought prompting introduces intermediate reasoning steps by encouraging the model to articulate its thought process before reaching a conclusion. Instead of predicting the final answer directly, the prompt guides the model to perform step-by-step reasoning. This has been shown to improve multi-step reasoning performance on complex tasks (Wei et al., 2022).
Logic-LM.
Logic-LM converts natural language statements into symbolic logic representations and uses a rule-based logical engine to carry out reasoning. By explicitly leveraging classical FOL rules, it enhances reasoning accuracy and structure, especially in tasks requiring strict logical consistency (Pan et al., 2023).
SymbCoT.
SymbCoT extends CoT prompting by incorporating symbolic representations and logic-based rules into the reasoning framework. This innovative approach first transforms natural language input into structured symbolic expressions, creating a FOL foundation for subsequent reasoning steps. The system then prompts the language model to reason over these symbolic representations with the assistance of FOL constraints and rules. This integration fundamentally improves the model’s capability for structured and formally grounded reasoning by bridging the gap between natural language understanding and symbolic logic (Xu et al., 2024b).
Aristotle.
Aristotle is a logic-complete reasoning framework designed for complex logical tasks by integrating symbolic structures across all reasoning stages. It includes three main components: Logical Decomposer, Logical Search Router, and Logical Resolver. These components enable structured task decomposition, guided reasoning, and contradiction resolution. Aristotle is particularly effective when used with LLMs, which benefit from symbolic supervision and structured guidance, leading to stronger performance than smaller models or purely generative strategies (Xu et al., 2024a).
@@ D.1 Statistics
Table 5: Basic Statistics of RepublicQA Benchmark
Metric Value
Total samples 200
Unique topics 58
Avg. context length 637.42
Avg. proposition length 166.42
Avg. explanation steps 6.10
Explanation completeness 100.0%
Basic Statistics. The RepublicQA benchmark comprises 200 carefully constructed samples covering 58 unique philosophical topics. Table 5 presents the fundamental statistical characteristics of the benchmark.
@@ D.2 Philosophical Concepts
The RepublicQA benchmark is rich in philosophical content, drawing directly from the core themes of Plato’s Republic. An analysis of the frequency and interconnections of key philosophical concepts provides insight into the benchmark’s thematic depth. Figure 8(a) visualizes the frequency of these concepts.
As shown in the Figure 8(a), Justice is the most prevalent concept, appearing 1308 times, revealing its central role in the Republic. Other frequently occurring concepts include State (846 occurrences), Soul (670 occurrences), Art (451 occurrences), Knowledge (242 occurrences), Virtue (206 occurrences), and Education (117 occurrences). The frequent occurences of Justice and State reflect Plato’s exploration of the just city and the just individual.
@@ D.3 Topic Modeling Results
D.4 Complexity Analysis of RepublicQA
RepublicQA presents a uniquely challenging benchmark for reasoning models due to its rich semantic diversity and structural complexity. Topic modeling reveals five core clusters corresponding to key themes in Platonic philosophy, ensuring evaluation across varied reasoning contexts rather than a single domain. This semantic distribution requires models to integrate multiple philosophical concepts simultaneously, making superficial pattern matching insufficient. In addition, the benchmark exhibits a balanced distribution of answer types and notable linguistic complexity, fostering evaluation scenarios that demand abstraction, disambiguation, and compositional reasoning.
We assess the linguistic and semantic complexity of RepublicQA by comparing it with four standard reasoning benchmarks: ProntoQA, FOLIO, ProofWriter, and ProverQA. Our analysis covers both surface and structural attributes using established metrics.
Linguistic Complexity. We compute the Flesch-Kincaid Grade Level (FKGL) to measure syntactic difficulty:
$\text{FKGL}=0.39\cdot\frac{N_{\text{words}}}{N_{\text{sentences}}}+11.8\cdot\frac{N_{\text{syllables}}}{N_{\text{words}}}-15.59$
Higher FKGL scores indicate more complex sentence structures and denser vocabulary.
Semantic Diversity. We report Type-Token Ratio (TTR) for vocabulary richness, MTLD for sustained lexical variety, and Unique Bigrams Ratio (UBR) for phrasal diversity. These metrics capture both abstraction and expression variability.
@@ D.4 Complexity Analysis of RepublicQA
Table 6: Complexity Metrics Comparison (Higher is Better)
Benchmark FKGL $\uparrow$ TTR $\uparrow$ MTLD $\uparrow$ UBR $\uparrow$ Vocab
ProntoQA (Saparov & He, 2022) 6.78 0.448 13.93 0.852 69
ProofWriter (Tafjord et al., 2020) 1.25 0.193 11.31 0.513 61
FOLIO (Han et al., 2022) 6.62 0.569 33.54 0.805 1,021
ProverQA (Qi et al., 2025) 8.41 0.362 49.43 0.639 4,891
RepublicQA (Ours) 11.94 0.685 74.81 0.929 2,083
@@ Appendix E Case Study
E.1 Dual-Form Representation and Semantic Precision
LogicAgent processes each proposition using both natural language and FOL representations. This dual-form design retains the conceptual richness of natural language while enabling symbolic reasoning under FOL. In the selected case, natural language captures nuanced distinctions (e.g., “justice” vs. “ability”), while FOL clarifies logical scope and reasoning structure. This integration supports:
@@ E.1 Dual-Form Representation and Semantic Precision
E.2 Multi-Perspective Reasoning for Robust Evaluation
@@
E.3 Planning, Execution, and Reflective Correction
@@
This case demonstrates how LogicAgent’s integrated reasoning pipeline, which combines dual-form representation, multi-perspective structure, and reflective correction, enables robust and explainable reasoning in semantically rich domains.
Appendix F Error Analysis
@@
Benchmark Semantic Richness Impact. As shown in Figure 10, the proportion of contrary relations varies widely across benchmarks. RepublicQA contains the richest semantic oppositions, with about 73% of instances involving meaningful contraries (e.g., justice vs. injustice). In contrast, FOLIO and ProverQA fall below 50%, reflecting limited availability of contraries despite their complexity. ProntoQA and ProofWriter contain almost none, restricting reasoning largely to contradictory checks. Notably, even in RepublicQA, certain propositions cannot yield valid contraries, underscoring that constructing oppositions is inherently difficult due to semantic ambiguity and hidden contextual links. Moreover, Figure 10 shows that accuracy on contrary cases is consistently lower than the overall accuracy, highlighting that contrary reasoning is intrinsically harder and more error-prone.
@@
Planning Limitations. Our framework does not strengthen the base model’s inherent planning ability. When reasoning paths are poorly identified, semantic analysis alone cannot recover performance, particularly on long-horizon datasets such as ProverQA. Moreover, the planner may inadvertently lengthen reasoning trajectories: ProverQA targets 6–9 step solutions, yet our full framework often extends beyond 10 steps, yielding an accuracy of 68.6%. In contrast, removing planning (woPlan) keeps reasoning within the expected range (6–9 steps) and achieves higher accuracy of 75% (Table 3). This suggests that reasoning length has a critical threshold, beyond which accuracy deteriorates.
@@
F.1 Computational Efficiency Analysis
@@
Time-Accuracy Trade-offs. As shown in Figure 5(b), LogicAgent exhibits a clear trade-off between reasoning accuracy and computational efficiency. Incorporating planning increases per-sample processing time by approximately 78%, but this additional cost yields structured, goal-oriented reasoning trajectories that are essential for tackling complex multi-hop logic tasks. Similarly, FOL translation contributes significantly to reasoning accuracy by enabling symbolic deduction and FOL consistency checking, yet it introduces substantial latency due to its multi-step transformation and verification pipeline. In contrast, relying solely on natural language input offers faster reasoning with reduced cost but sacrifices FOL rigor and precision in logical reasoning. These results underscore the value of structured symbolic reasoning, particularly when accuracy and interpretability are prioritized.
Table 7: Token Consumption Analysis for RepublicQA
@@
Stage-Level Timing Breakdown.
To understand intra-system efficiency, we analyze stage-wise processing time for LogicAgent’s full configuration on RepublicQA (Table 8). The Logical Reasoning Stage is the dominant computational bottleneck, accounting for 75.1% of total runtime (139.66s). This is attributed to three interacting factors: (1) multi-path execution across $S_{1}$ and $\lnot S_{1}$, effectively doubling reasoning steps; (2) context-to-FOL translation involving semantic disambiguation and quantifier binding; and (3) execution of detailed plans that enforce step-wise logical progression. The Semantic Structuring Stage, responsible for constructing Greimas’ semiotic square, is comparatively efficient (13.6%), as it involves short-form outputs and deterministic linguistic transformations. The Reflective Verification Stage, while occasionally expensive, benefits from its adaptive design. In most cases, it executes lightweight verification; only uncertain or conflicting cases invoke full re-analysis, keeping its average cost low (11.3%).
@@
Appendix G Full Prompting

### v5

@@ B.1 Formal Syntax and Validation of FOL
FOL CFG Grammar. To ensure the well-formedness of FOL expressions, we implement a symbolic parser using the nltk library (Bird, 2006). Specifically, we define a context-free grammar (CFG) to support automatic parsing and validation of logical formulas throughout our pipeline:
@@ C.1 Benchmarks
RepublicQA is a philosophical reasoning benchmark derived from classical works in Western philosophy, including Plato’s Republic (Plato, 2016), Aristotle’s Metaphysics (Aristoteles and Apostle, 1966), and the Nicomachean Ethics (Irwin and others, 2019). These traditions provide rich discussions of justice, morality, governance, virtue, and knowledge, yielding abstract propositions and structured counterarguments that are well suited for evaluating advanced reasoning. The benchmark presents complex logical problems in which models must judge whether philosophical statements follow from contextual premises. RepublicQA assesses the ability to engage with abstract concepts, moral and ethical reasoning, and classical argumentative patterns while preserving high complexity in both language and inference. Each example reflects foundational questions in Western philosophy and requires reasoning over conditional claims, normative principles, and abstract conceptual relations. To maintain logical consistency, the examples are organized around classical philosophical dialogues and argumentative structures that require multi-step reasoning to determine whether conclusions about justice, virtue, or political order are supported by the given premises.
@@ C.2 Baselines
Naive Prompting. The model directly receives the question and produces an answer without guidance or intermediate steps. Reasoning is neither encouraged nor structured, making this approach suitable only for simple factual queries.
Chain-of-Thought (CoT). CoT prompting elicits step-by-step reasoning before the final answer, improving multi-step reasoning performance by explicitly revealing intermediate steps (Wei et al., 2022).
Cumulative Reasoning (CR). CR iteratively refines reasoning across multiple passes. Intermediate outputs from earlier steps serve as inputs to later ones, enabling gradual accumulation and refinement of reasoning (Zhang et al., 2023).
Tree-of-Thought (ToT). ToT explores reasoning as a search tree. Instead of a single chain, multiple reasoning paths are generated, evaluated, and pruned, allowing the model to retain only the most promising trajectories (Yao et al., 2023).
Logic-LM. Logic-LM translates natural language into first-order logic and applies symbolic solvers for rule-based deduction. This enhances structure and consistency, especially for tasks requiring strict logical validity (Pan et al., 2023).
SymbCoT. SymbCoT augments CoT with symbolic representations and logic constraints. Natural language inputs are converted into symbolic forms, and reasoning proceeds under formal logical guidance (Xu et al., 2024b).
Aristotle. Aristotle is a logic-complete framework integrating symbolic structures throughout the reasoning pipeline. Its Logical Decomposer, Logical Search Router, and Logical Resolver support structured decomposition, guided search, and contradiction handling, enabling strong performance on complex logical tasks (Xu et al., 2024a).
Figure 7: Answer distribution across different benchmarks.
Table 6: Benchmark Comparison: Basic Statistics and Semantic Complexity. Bold indicates the best performance and underline indicates the second best.
Basic Statistics Semantic Complexity
[…]
ProntoQA 500 1 69 11 6.78 0.448 13.93 0.852 0.00
FOLIO 204 1 1,021 0* 6.62 0.569 33.54 0.805 0.30
ProofWriter 600 345 61 0* 1.25 0.193 11.31 0.513 0.00
ProverQA 500 500 2,453 25.73 8.44 0.616 34.84 0.774 0.13
RepublicQA 600 61 4,070 16.22 11.94 0.685 74.81 0.929 0.70
+41.5% +11.2% +114.7% +9.0% +133.3%
*Dataset does not provide explicit reasoning steps.
@@ D.1 Statistics
Basic Statistics. The RepublicQA benchmark comprises 600 carefully constructed samples covering 61 unique philosophical topics. Table 6 presents the fundamental statistical characteristics of the benchmark.
@@ D.2 Philosophical Concepts
RepublicQA is deeply grounded in the thematic structure of Plato’s Republic, and its philosophical concepts directly shape both the semantic and logical complexity of the benchmark. As shown in Figure 8(a), core concepts such as Justice (1,308 occurrences), State (846), Soul (670), Art (451), Knowledge (242), Virtue (206), and Education (117) appear frequently throughout the dataset. These abstract and interrelated notions introduce substantial semantic richness and require models to integrate multiple conceptual layers when drawing conclusions. Their interactions create reasoning scenarios that are considerably deeper and more context dependent than those found in benchmarks built around concrete entities or isolated factual predicates.
@@ D.3 Topic Modeling Results
D.4 Semantic Complexity Analysis
To assess the semantic complexity of RepublicQA relative to existing reasoning benchmarks, we evaluate all datasets across three complementary dimensions, as summarized in Table 6: (1) conceptual complexity, measured by FKGL; (2) lexical diversity, captured by TTR, MTLD, and UBR; and (3) structural contrast, quantified through the Contrary metric.
Conceptual Complexity. We measure sentence-level abstraction using the Flesch–Kincaid Grade Level (FKGL):
$\text{FKGL}=0.39\cdot\frac{N_{\text{words}}}{N_{\text{sentences}}}+11.8\cdot\frac{N_{\text{syllables}}}{N_{\text{words}}}-15.59.$
Higher scores indicate denser conceptual content and more syntactically demanding propositions.
Lexical Diversity. We assess vocabulary and phrasal variation through three complementary measures: Type–Token Ratio (TTR), MTLD for long-span lexical variety, and Unique Bigrams Ratio (UBR) for phrasal diversity. These metrics capture the breadth and stability of semantic expression beyond surface-level repetition.
Structural Contrast. To quantify higher-level semantic structure, we use the Contrary metric, which measures the presence of systematically constructed contrasting relations within each dataset. Higher values correspond to richer semantic tension and more nuanced relational patterns that require models to integrate multiple, potentially competing interpretations.
Results. As shown in Table 6, RepublicQA substantially surpasses existing benchmarks across all dimensions. It requires college-level reading (FKGL = 11.94), exhibits markedly richer lexical diversity (TTR = 0.685), maintains long-span expressive variability (MTLD = 74.81), and achieves a high phrasal diversity (UBR = 0.929). In addition, RepublicQA uniquely incorporates systematically constructed contrary relations (Contrary = 0.70), introducing semantic tension and multi-perspective reasoning that are absent from rule-based datasets such as ProntoQA and ProofWriter.
@@ D.4 Semantic Complexity Analysis
Dataset Simple Medium Hard
RepublicQA NL: Debt repayment is a moral obligation.
FOL: $\forall a\forall x(\mathrm{Repay}(a,x)\rightarrow\mathrm{MoralObligation}(a))$
NL: Some debts originate from unjust or fraudulent means.
FOL: $\exists x(\mathrm{Debt}(x)\land(\mathrm{Fraudulent}(x)\lor\mathrm{Unjust}(x)))$
NL: Justice requires helping friends and avoiding harm to innocents.
FOL: $\forall a(\mathrm{Just}(a)\rightarrow(\forall y(\mathrm{Friend}(y)\rightarrow\mathrm{Beneficial}(a,y))\land\forall z(\mathrm{Innocent}(z)\rightarrow\lnot\mathrm{Harm}(a,z))))$
ProverQA NL: Loyal is well-trained.
FOL: $\mathrm{WellTrained}(\mathit{loyal})$
NL: If Legend has strong hooves and a powerful gait, he can be a champion.
FOL: $(\mathrm{StrongHooves}(\ell)\land\mathrm{PowerfulGait}(\ell))\rightarrow\mathrm{CanBeChampion}(\ell)$
NL: If Legend is competitive, then he has (unique color xor distinctive marking), is good-tempered, not athletic, etc.
FOL: $\mathrm{Competitive}(\ell)\rightarrow((\mathrm{UniqueColor}(\ell)\oplus\mathrm{DistinctiveMarking}(\ell))\land\mathrm{GoodTemperament}(\ell)\land\lnot\mathrm{AthleticBuild}(\ell)\dots)$
ProofWriter NL: Charlie is kind.
FOL: $\mathrm{Kind}(\mathit{charlie},\mathrm{True})$
NL: If someone is quiet and cold, they are smart.
FOL: $\forall x((\mathrm{Quiet}(x)\land\mathrm{Cold}(x))\rightarrow\mathrm{Smart}(x))$
NL: Rough is Cold; Cold and Smart is Red; Red is Rough (cyclic reasoning chain).
FOL: $\mathrm{Rough}(x)\rightarrow\mathrm{Cold}(x);\ (\mathrm{Cold}(x)\land\mathrm{Smart}(x))\rightarrow\mathrm{Red}(x);\ \mathrm{Red}(x)\rightarrow\mathrm{Rough}(x)$
FOLIO NL: If people perform, they attend school events.
FOL: $\forall x(\mathrm{Perform}(x)\rightarrow\mathrm{AttendEngage}(x))$
NL: Inactive people chaperone school dances.
FOL: $\forall x(\mathrm{InactiveDisinterested}(x)\rightarrow\mathrm{ChaperoneDances}(x))$
NL: Bonnie either attends events as a student, or neither.
FOL: $\mathrm{AttendEngage}(\mathit{bonnie})\land\mathrm{StudentSchool}(\mathit{bonnie})\ \lor\ \lnot(\mathrm{AttendEngage}(\mathit{bonnie})\land\mathrm{StudentSchool}(\mathit{bonnie}))$
@@ Appendix E Case Study
Dual-Form Representation and Semantic Precision
LogicAgent processes each proposition using both natural language and FOL representations. This dual-form design retains the conceptual richness of natural language while enabling symbolic reasoning under FOL. In the selected case, natural language captures nuanced distinctions (e.g., “justice” vs. “ability”), while FOL clarifies logical scope and reasoning structure:
@@
[…]
Multi-Perspective Reasoning for Robust Evaluation
@@
Planning, Execution, and Reflective Correction
@@ Appendix F Error Analysis
Benchmark Semantic Richness Impact. As shown in Figure 10, the availability of valid contraries differs substantially across benchmarks. RepublicQA exhibits the richest set of meaningful conceptual contrasts, while FOLIO, ProverQA, ProntoQA, and ProofWriter contain far fewer, which limits opportunities for multi-perspective reasoning. Even within RepublicQA, not all propositions admit well-defined contraries, since constructing them requires resolving semantic ambiguity, aligning abstract concepts, and recovering context-dependent links that are often implicit. Figure 10 further shows that accuracy on contrary cases is consistently lower than overall accuracy, indicating that contrary reasoning remains intrinsically difficult for current models.
@@
Planning Limitations. Our framework does not enhance the base model’s intrinsic planning ability. When reasoning paths are poorly estimated, semantic analysis alone cannot compensate, especially on long-horizon datasets such as ProverQA. The planner may also over-extend reasoning: while ProverQA typically requires 6–9 steps, our full model often exceeds 10 steps and drops to 68.6% accuracy. In contrast, removing planning (woPlan) keeps trajectories within the expected range and achieves 75% accuracy (Table 3). These results indicate that reasoning length exhibits a critical threshold beyond which accuracy degrades sharply.
@@
Appendix G Computational Efficiency Analysis
@@
Time-Accuracy Trade-offs. As shown in Figure 5(b), LogicAgent demonstrates a clear trade-off between accuracy and efficiency. Planning increases computation time by roughly 78% but provides structured, goal-directed trajectories that benefit complex multi-hop reasoning. FOL translation further boosts accuracy through symbolic deduction and consistency checking, though at the cost of substantial latency. In contrast, purely natural language reasoning is faster but lacks the rigor and precision afforded by symbolic structure. These findings highlight the importance of structured reasoning when accuracy and interpretability are required.
Table 8: Token Consumption Analysis
@@
Stage-Level Timing Breakdown. To understand intra-system efficiency, we analyze stage-wise processing time for LogicAgent’s full configuration on RepublicQA (Table 9). The Logical Reasoning Stage is the dominant computational bottleneck, accounting for 75.1% of total runtime (139.66s). This is attributed to three interacting factors: (1) multi-path execution across $S_{1}$ and $\lnot S_{1}$, effectively doubling reasoning steps; (2) context-to-FOL translation involving semantic disambiguation and quantifier binding; and (3) execution of detailed plans that enforce step-wise logical progression. The Semantic Structuring Stage, responsible for constructing Greimas’ semiotic square, is comparatively efficient (13.6%), as it involves short-form outputs and deterministic linguistic transformations. The Reflective Verification Stage, while occasionally expensive, benefits from its adaptive design. In most cases, it executes lightweight verification; only uncertain or conflicting cases invoke full re-analysis, keeping its average cost low (11.3%).
@@
Appendix H Full Prompting